sum of first two term f a decreasing g.p is 5/4 and sum of infinite term is 9/4. find common ratio and first term

Grade:12th pass

1 Answers

Shubham Raj
7 years ago
Let a be first term and r be common ratio.a+ar=5/4a/1-r =9/4Solve both equations to get a=3/4 ,r=2/3